Tartaric acid (C4H6O6) is naturally found in fruits like grapes and banana. It is used in the food industry as a leaving agent that gives a sour taste. Determine the percent composition of each element in the compound.

1
Expert's answer
2022-02-11T11:00:03-0500

Using RAM of C=12, H=1, and O=16

RFM of C4H6O6 = (12x4)+(1x6)+(16x6)

=48+6+96 = 150g/mol


% by mass of C = (48/150)x100 = 32%


% by mass of H = (6/150)x100 = 4%


% by mass of O = (96/150)x100 = 64%